# Performance-lossless Black-box Model Watermarking

### v1

@@
Abstract.
With the development of deep learning, high-value and high-cost models have become valuable assets, and related intellectual property protection technologies have become a hot topic. However, existing model watermarking work in black-box scenarios mainly originates from training-based backdoor methods, which probably degrade original task performance. To address this, we propose a branch backdoor-based model watermarking protocol to protect model intellectual property, where a construction based on a message authentication scheme is adopted as the branch indicator. We prove the lossless performance of the protocol by reduction. Taking the language generation task as an instance, we show the effectiveness of the proposed protocol.
performance-lossless; black-box; watermark; branch
1. Introduction
Recently, the development of deep learning technology represented by GPTs (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023b), GANs (Goodfellow et al., 2014; Brock et al., 2019; Karras et al., 2021), Diffusion models (Dhariwal and Nichol, 2021; Ho et al., 2022; Rombach et al., 2022), etc. has become increasingly mature. Applications (OpenAI, 2023a; Google, 2023; Anthropic, 2023) in Computer Vision (CV), Natural Language Processing (NLP), etc., have demonstrated tremendous commercial value. However, developing high-performance models entails substantial costs, including the involvement of machine learning experts with rich experience, extensive and high-quality task datasets, and expensive computational resources. Development teams maintain legitimate intellectual property rights over their produced models and often refrain from open-sourcing critical model details. They provide on-demand model functionality services to users through a commercial paradigm of remote access to model input-output APIs, such as OpenAI’s ChatGPT (OpenAI, 2023a), Google’s Bard (Google, 2023), and Anthropic’s Claude (Anthropic, 2023).
@@ 1. Introduction
Watermarking serves as an effective solution for protecting intellectual property. Existing model watermark works can generally be divided into three types based on the model knowledge needed for copyright verification or the ability to control the model: white-box watermarks, black-box watermarks, and no-box watermarks. White-box verification watermarking assumes that the model is completely transparent, allowing the verifier to manipulate the model within the verification rules to detect the watermark signal. In black-box verification watermarking scenarios, the verifier lacks knowledge of the inner workings of the model, but can manipulate the inputs of the model to obtain corresponding outputs to detect the watermark signal. In the case of no-box validation watermarks, the verifier only has access to the model’s output to detect the watermark signal. In the threat scenario described previously, suspect models are difficult to obtain, which limits white-box model watermarking methods. No-box model watermarking requires independently meaningful model outputs, which makes it useful only for generative tasks. Therefore, black-box model watermarking is more common and practical.
@@
Black-box model watermarking works mainly rely on backdoor techniques. In 2018, Adi et al. (Adi et al., 2018) first established the connection between backdoor models and model watermarks, introducing the first backdoor-based model watermark work. Verifiers activate a pre-embedded backdoor in the model by inputting trigger samples during the inference process and detecting watermark signals in the output for model ownership confirmation. Existing backdoor-based model watermarks (Darvish Rouhani et al., 2019; Le Merrer et al., 2020; Adi et al., 2018; Chen et al., 2019; Li et al., 2022b; Jia et al., 2022; Lounici et al., 2021; Lao et al., 2022a; Jia et al., 2021; Guo and Potkonjak, 2018; Zhu et al., 2020; Maung Maung and Kiya, 2021; Li et al., 2022a, 2019; Lao et al., 2022b; Yang et al., 2021; Charette et al., 2022; Namba and Sakuma, 2019; Bansal et al., 2022; Zhong et al., 2020; Mehta et al., 2022) usually involve fine-tuning or retraining the model’s weights to embed the backdoor, which essentially constitutes multitask learning. Due to the incomprehensible nature of neural networks, despite attempts to separate the influences between the two tasks through carefully designed loss functions (Jia et al., 2021), weight selections (Lao et al., 2022a; Namba and Sakuma, 2019; Bansal et al., 2022), trigger sets (Darvish Rouhani et al., 2019; Adi et al., 2018; Zhang et al., 2018; Jia et al., 2022; Lounici et al., 2021; Lao et al., 2022a; Jia et al., 2021; Guo and Potkonjak, 2018; Zhu et al., 2020; Maung Maung and Kiya, 2021; Zhong et al., 2020), etc., there is no reliable theoretical analysis to prove that the watermarking task will not distort the normal task. Their experimental results also indicate that embedding watermarks impacts the accuracy of image classification models.
An ideal model watermark should not affect the model’s original functionality. In order to meet this requirement, we propose a performance-lossless black-box model watermarking, with specific contributions as follows:
  1. (1)
    We propose a branch watermarking protocol for black-box model copyright verification scenarios, which is a watermark embedding paradigm during the model service deployment phase. The protocol decouples the original task of the model from the verification task by inserting an independent watermark module in parallel to the target model within the API.
  2. (2)
    To achieve the watermark module’s independence, we propose the concept of the unlearnable trigger based on the provably unforgeable Message Authentication Code (MAC), and our constructions enjoy a formal security analysis that reduces the performance-lossless property of the model watermarking to the properties of the underlying cryptographic primitive.
  3. (3)
    Taking the language generation model as an example, we provide the instantiation strategy of the proposed watermark protocol, and combine it with the existing no-box watermarking scheme to improve the security against interference attacks.
2. Related Work
2.1. Backdoor Attacks
@@
Backdoor attacks aim to embed hidden backdoors into deep neural networks, allowing the model to perform normally on regular samples but activate the internal backdoor when exposed to predefined triggers, resulting in malicious or poor-quality predictions. Depending on whether backdoor implantation involves training a victim model, these attacks are mainly divided into two types: training-based backdoors and training-free backdoors.
2.1.1. Training-based Backdoors
We can think of the initial training phase of the model and the fine-tuning phase of subsequent downstream tasks as the development phase of the model. Most backdoor attacks that occur during the development phase fall into the training-based backdoor category, whether by poisoning the training dataset or injecting malicious code into the training code.
@@ 2.1.1. Training-based Backdoors
Backdoor attacks based on poisoned data (Gu et al., 2017; Turner et al., 2019; Chen et al., 2017; Saha et al., 2020; Turner et al., 2018) have long been a topic of concern. When model developers face resource constraints, whether due to a lack of training data or limited computing power, they may purchase datasets from third parties or delegate training tasks to third parties, thereby exposing themselves to the risk of attackers injecting poisoned data into the training data. Attackers poison the training data, causing the model to learn the knowledge of the poisoned data during the training process to plant the backdoor. Backdoor attacks based on poisoned training code (Bagdasaryan and Shmatikov, 2021) constitute a form of blind attack on unknown models. Attackers insert malicious code into the model’s training code, automatically poisons the data and controls the loss function, and injects new tasks into the model during the training process to plant backdoors.
@@
The knowledge associated with these poisoned data often leads to poor predictions. However, this backdoor technique can also be used for model copyright verification when the knowledge is used as a signature of model developers.
2.1.2. Training-free Backdoors
Compared with training-based backdoor methods, training-free backdoor methods inject backdoors completely independently of the training process of the model. Attackers insert backdoors by injecting poisonous branches into the model’s inference process during the model deployment phase, such as poisoning model structures or control flow. This type of method does not require secondary modification of the model weights, thereby avoiding damaging the original performance of the model.
Backdoor attacks based on the poisoned model structure (Liu et al., 2022; Zou et al., 2018; Tang et al., 2020; Bober-Irizar et al., 2023) involve attackers adding additional subnetworks to the victim model’s network structure to inject a backdoor. It is worth noting that the poisoned structure is only used during the inference process and will not affect the training process. Backdoor attacks based on poisoned control flow (Li et al., 2021; Clifford et al., 2022; Wang, 2022) occur when the victim model is deployed and run as part of software. Attackers insert new node modules into the control flow to inject a backdoor. Both attack types essentially involve inserting a branch into the victim model’s inference process, with the difference lying in the form of the backdoor trigger detector, whether it is a neural network or a diverse structure.
The mentioned attacks, particularly control flow poisoning, resemble traditional software attack techniques such as Trojans. The principles for designing an ideal Trojan horse attack, as proposed in the work (Tang et al., 2020), provide guidance for designing our branch watermarking protocol: 1) The Trojan horse attack should be model-agnostic, meaning it can be applied to different DNNs with minimal effort; 2) Inserting a Trojan horse in the model should not change the performance on the original task; 3) The triggers should be diverse; 4) The Trojan horse should be invisible and not detectable by existing Trojan horse detection algorithms.
2.2. Watermarking for Models
@@
Model watermarking is a technology used to protect and track machine learning models, similar to the application of digital watermarks in digital media. The main goal of model watermarking is to embed a unique identifier or label in the model without affecting the performance or accuracy of the model, thereby identifying and verifying the legitimacy of the model. This helps prevent unauthorized copying, misuse or theft of the model while also serving as a means of proving the provenance and integrity of the model. In the context of the remote serving paradigm of models, we mainly focus on two types of model watermarks: black-box watermarks and no-box watermarks.
2.2.1. Black-box Watermarking
@@
The “black box” is an image description of the verifier’s knowledge of the model, indicating that the verifier’s knowledge of the model is limited to queries and responses. Model owners use techniques such as backdoor attacks to embed signature-related watermark information in the model. They use specific trigger query-response pairs to detect watermarks for verification.
@@
The mentioned Adi et al.’s work (Adi et al., 2018) introduced a relationship between malicious backdoor attacks and benign watermark defenses for the model, reducing the watermark task of deep neural networks to a backdoor task. Existing model watermarking approaches based on backdoor design share the same thought as methods such as training-based backdoors. Model owners use specific input-output pairs as watermark patterns, which is a form of knowledge that needs to be learned. This knowledge is injected into the model and trained to create backdoors as watermarks. Then, during ownership disputes, specific input triggers are used to activate the watermark backdoor and obtain relevant specific outputs for copyright verification. However, as discussed in section 2.1.2, these backdoor attacks rely on injecting the backdoor during training. The backdoor task and the model’s original task are entangled during the training process, which makes it challenging to ensure that the original task performance is not distorted. The same goes for watermark tasks.
2.2.2. No-box Watermarking
@@
For tasks related to copyright protection of language generation models, methods such as synonym replacement-based no-box watermarking were proposed, and subsequent work further improved the watermark scheme’s impact on text quality (He et al., 2022b, a). Kirchenbauer et al. (Kirchenbauer et al., 2023) proposed a zero-bit watermark scheme that adds biased signals based on blacklists and whitelists to the outputs (i.e., logits) of the model. While the paper targeted a binary classification traceability task for human-machine text, it can be applied to model copyright verification tasks in a no-box scenario. Some works (Lee et al., 2023; Zhao et al., 2023a; Fu et al., 2023; Liu et al., 2023; Yoo et al., 2023) have extended and improved upon this work (Kirchenbauer et al., 2023). Zhao et al. (Zhao et al., 2023b) transformed the model watermark work (Charette et al., 2022) designed for image classification tasks, to language generation tasks in the context of no-box watermarking.
3. A Performance-lossless Branch Watermarking Protocol for Model
In this section, we provide a comprehensive description of the threat model targeted in this paper. We will then propose the corresponding model watermarking protocol and analyze the potential attacks that the watermarking protocol may face. Finally, we summarize the threat models and attack methods and outline the principles that watermarking protocols should satisfy in practical applications.
Figure 1. There are three roles in the threat scenario. The model development team provide black-box deployment to the application service provider. The latter may dishonestly provide application services to end users, which we summarize as two cases of dishonesty in the red zone.
3.1. Threat Model
@@
The application service provider is a team of non-machine learning experts who purchase model functionality services from the model owner. They leverage remote APIs to deploy specific applications (such as chatbots) and deliver well-encapsulated, user-friendly, and easy-to-use model applications to end-users.
@@
  2. (2)
    The application provider uses a well-known high-performance model A through the remote API but deliberately claims that it is a self-developed model B. This allows them to gain market share and profits while compromising the intellectual property of model owners.
@@
3.2. Planting A Parallel Branch in Model API
Remote services have the advantages of a black box model, allowing watermark modules to be added to the API. We propose a branch backdoor-based watermarking protocol for model. The idea of the mutual transformation between malicious backdoors and benign watermarks initially originated from Adi et al.’s work (Adi et al., 2018), but unlike previous work, we use a training-free backdoor that modifies the control flow rather than a training-based backdoor. The advantage of this is that it can decouple the original tasks and verification tasks.
3.2.1. Components
@@
The protocol consists of two modules: the model and the watermark, with the watermark module further divided into the inner module and the outer module. The outer module includes a trigger generator and evidence verifier submodules. The inner module includes a trigger detector and evidence generator submodules deployed within the API.
3.2.2. Overview
@@
Figure 2 provides a complete process of the branch watermarking protocol for the model. When an ownership dispute arises, the verifier uses the outer module to generate triggers, sends these triggers as query requests to the API, and verifies the returned responses to complete copyright certification. Specifically, when the API receives a query request, it forwards the input content to both the model module and the inner module, initially running in the default service state. The model module performs the original inference process to obtain the output content.
@@
Meanwhile, the inner module checks whether the current input is a trigger. If it is, the API switches to the forensic state, executing the evidence generator module to inject ownership information into the model’s original output content, generating ownership evidence, which the API returns as a response to the user. If the input is not a trigger, the API continues in the service state, returning the model’s original output as a response to the user. Importantly, the outer module is private or only open to a trusted third party.
@@
Faced with the two situations of dishonest behavior in the threat model, the owner of model A can naturally use the outer module to generate triggers, conduct investigations on suspicious remote application services, and verify the evidence. In the first case, sending triggers to the suspicious application that pretends to use model A would yield no corresponding ownership evidence. In the second case, sending triggers to an application claiming to use model B but with performance closely resembling model A would yield corresponding ownership evidence.
Figure 2. There are two modules in the IP verification protocol. The outer module is deployed outside the model API, responsible for generating triggers to obtain model copyright evidence and verifying the evidence. The inner module is added as a branch in the model API, responsible for checking whether the input is a trigger and generating the copyright evidence.
3.2.3. Defining Watermarking Protocol

### v2

@@
Abstract
With the development of deep learning, high-value and high-cost models have become valuable assets, and related intellectual property protection technologies have become a hot topic. However, existing model watermarking work in black-box scenarios mainly originates from training-based backdoor methods, which probably degrade primary task performance. To address this, we propose a branch backdoor-based model watermarking protocol to protect model intellectual property, where a construction based on a message authentication scheme is adopted as the branch indicator after a comparative analysis with secure cryptographic technologies primitives. We prove the lossless performance of the protocol by reduction. In addition, we analyze the potential threats to the protocol and provide a secure and feasible watermarking instance for language models.
Index Terms:
Performance-lossless, black-box, watermark, branch.
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
Recently, the development of deep learning technology, represented by GPTs [1, 2, 3, 4], GANs [5, 6, 7], Diffusion models [8, 9, 10], etc., has become increasingly mature. Its applications [11, 12, 13] in Computer Vision (CV), Natural Language Processing (NLP), etc., have demonstrated tremendous commercial value. However, developing high-performance models entails substantial costs, including the involvement of machine learning experts with rich experience, extensive and high-quality task datasets, and expensive computational resources. Development teams maintain legitimate intellectual property rights over their produced models and often refrain from open-sourcing critical model details. They provide on-demand model functionality services to users through a commercial paradigm of remote access to model input-output Application Programming Interfaces (APIs), such as OpenAI’s ChatGPT [11], Google’s Bard [12], and Anthropic’s Claude [13].
@@ I Introduction
Watermarking serves as an effective solution for protecting intellectual property. Existing model watermarking works can generally be divided into three types based on the model knowledge needed for copyright verification or the ability to control the model: white-box watermarking, black-box watermarking, and no-box watermarking. White-box verification watermarking assumes that the model is completely transparent, allowing the verifier to manipulate the model within the verification rules to detect the watermark signal. In black-box verification watermarking scenarios, the verifier lacks knowledge of the inner workings of the model but can manipulate the inputs of the model to obtain specific outputs, thereby detecting the watermark signal. In the case of no-box validation watermarking, the verifier only has access to the model’s output to detect the watermark signal. In the black-box threat scenario, suspect models are difficult to obtain, which limits white-box model watermarking methods. No-box model watermarking requires independently meaningful model outputs, which makes it useful only for generative tasks. Therefore, black-box model watermarking is more practical.
@@
Black-box model watermarking works mainly rely on backdoor techniques. In 2018, Adi et al. [14] first established the connection between backdoor models and model watermarking, introducing the first backdoor-based model watermarking work. Verifiers activate a pre-embedded backdoor in the model by inputting trigger samples during the inference process and detecting watermark signals in the output for model ownership confirmation. Existing backdoor-based model watermarking works [15, 16, 14, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] usually involve fine-tuning or retraining the model’s weights to embed the backdoor, which essentially constitutes multitask learning. Due to the incomprehensible nature of neural networks, despite attempts to separate the influences between the two tasks through carefully designed loss functions [22], weight selections [21, 31, 32], trigger sets [15, 14, 35, 19, 20, 21, 22, 23, 24, 25, 33], etc., there is no reliable theoretical analysis to prove that the watermarking task will not distort the primary task. Their experimental results also indicate that embedding watermarks impacts the accuracy of image classification models.
An ideal model watermark should not affect the model’s original functionality. To meet this requirement, we adopt the idea proposed by Adi et al. [14] that watermarks can be transformed from backdoors and draw inspiration from the undetectable black-box construction in the work of Goldwasser et al. [36]. We propose a provably performance-lossless black-box model watermarking protocol named BranchWM. Specifically, we add a parallel branch for forensic tasks outside the original model structure, decoupling the original task of the model from the watermark forensic task. Then, we construct a trigger for forensic tasks using an existence unforgeable Message Authentication Code (MAC) scheme (a secure cryptographic primitive). Without owners’ keys and white-box access to the model, it is hard for attackers to know trigger patterns and forge triggers. Therefore, the probability that a normal sample triggers the forensic branch is computationally ignored, which leads to a performance-loss model watermarking. Finally, we provide an instantiation algorithm for our proposed watermark using language models and analyze potential interference attacks during actual forensic processes as well as improved watermark instances.
Our major contributions can be summarized as follows:
  1. 1.
    We first propose a provably performance-lossless black-box model watermarking protocol named BranchWM, which is based on the branch structure and secure cryptography.
  2. 2.
    We employ an MAC with existential unforgeability to construct a forensic trigger named Mac-based Unaccessible and Fast Trigger (MUFT), which is both unaccessible and can be quickly detected.
  3. 3.
    We analyze the potential attacks in the forensic process and provide a secure and feasible watermarking instance for language models.
II Related Work
II-A Backdoor Attacks
@@
Backdoor attacks aim to embed hidden backdoors into deep neural networks, allowing the model to perform normally on regular samples but activate the internal backdoor when exposed to predefined triggers, resulting in malicious or poor-quality predictions.
@@
Backdoor attacks based on poisoned data [37, 38, 39, 40, 41] have long been a topic of concern. When model developers face resource constraints, whether due to a lack of training data or limited computing power, they may purchase datasets from third parties or delegate training tasks to third parties, thereby exposing themselves to the risk of attackers injecting poisoned data into the training data. The attacker poisons the training data, causing the model to learn the knowledge of the poisoned data during the training process to plant the backdoor. Backdoor attacks based on poisoned training code [42] constitute a form of blind attack on unknown models. The attacker inserts malicious code into the model’s training code, automatically poisons the data and controls the loss function, and injects new tasks into the model during the training process to plant backdoors.
@@
Whether the work is based on poisoned data or poisoned training code, it usually involves injecting backdoors through training. The knowledge learned from this poisoned content often leads to poor predictions on specific triggers. When the knowledge is used as a signature of model developers, this backdoor technique can also be used for model copyright verification.
II-B Watermarking for Models
@@
Model watermarking is a technology used to protect and track machine learning models, similar to the application of digital watermarking in digital media. The main goal of model watermarking is to embed a unique identifier or label in the model without affecting the performance or accuracy of the model, thereby identifying and verifying the legitimacy of the model. This helps prevent unauthorized copying, misuse or theft of the model while also serving as a means of proving the provenance and integrity of the model. In the context of the remote serving paradigm of models, we mainly focus on two types of model watermarking: black-box watermarking and no-box watermarking.
II-B1 Black-box Watermarking
@@
The “black box” is an image description of the verifier’s knowledge of the model, meaning that the verifier’s knowledge of the model is limited to queries and responses. Model owners use techniques such as backdoor attacks to embed signature-related watermark information in the model. They use specific trigger query-response pairs to detect watermarks for verification.
@@
Adi et al. [14] introduced a relationship between malicious backdoor attacks and benign watermark defenses for the model, converting the watermarking task of deep neural networks to a backdoor task. Existing model watermarking approaches based on backdoor design share the same thought as methods such as training-based backdoors. Model owners use specific input-output pairs as watermark patterns, which is a form of knowledge that needs to be learned. This knowledge is injected by training the model. Then, during ownership disputes, specific input triggers are used to activate the watermark backdoor and obtain relevant specific outputs for copyright verification. However, these backdoor attacks rely on injecting the backdoor during training. The backdoor task and the model’s original task are entangled during the training process, which makes it challenging to ensure that the original task performance is not distorted. The same goes for watermarking tasks. Ensuring that the watermark knowledge does not impact the original task performance of the model is a problem that remains to be addressed theoretically.
II-B2 No-box Watermarking
@@
For tasks related to copyright protection of language generation models, methods such as synonym replacement-based no-box watermarking were proposed, and subsequent work further improved the impact of watermarking scheme on text quality [43, 44]. Kirchenbauer et al. [45] proposed a zero-bit watermarking scheme that adds biased signals based on greenlists and redlists to the outputs (i.e., logits) of the model. While the paper targeted a binary classification traceability task for human-machine text, it can be applied to model copyright verification tasks in a no-box scenario. Some works [46, 47, 48, 49, 50] have extended and improved upon this work [45]. Zhao et al. [51] transformed the model watermarking work [30] designed for image classification tasks to language generation tasks in the context of no-box watermarking. While no-box model watermarking is limited to generation tasks, it can be used to enhance the concealment of backdoor-based watermarking.
III A Performance-lossless Branch Watermarking Protocol for Model
In this section, we provide a comprehensive description of the threat model targeted in this paper. Then, we explain the motivation behind proposing a black-box performance-lossless branch backdoor-based watermarking for model. When introducing the proposed method, we first clarify the motivation and then provide a formal definition. We then present the core construction of the watermarking based on MAC and prove its performance-lossless property. Finally, to highlight the differences between our work and that of Goldwasser et al. [36], we analyze the impact of three secure cryptographic primitives——MAC, hash, and digital signatures—on the construction of performance-lossless branch watermarking and explain the advantages of using MAC.
Figure 1: There are three roles in the threat scenario. The model development team provide black-box deployment to the application service provider. The latter may dishonestly provide application services to end users, which we summarize as two cases of dishonesty in the red zone.
III-A Threat Model
@@
The application service provider is a team of non-machine learning experts who purchase model functionality services from the model owner. They leverage remote APIs to deploy specific applications (such as poe.com) and deliver well-encapsulated, user-friendly, and easy-to-use model applications to end-users.
@@
  2. 2.
    The application provider uses a well-known high-performance model A through the remote API but deliberately claims that it is self-developed model B. This allows them to gain market share and profits while compromising the intellectual property of model owners.
@@
III-B Motivation
Goldwasser et al. [36] defined the black-box undetectability of the backdoor model based on computational indistinguishability and constructed a black-box undetectable backdoor from a digital signature scheme that meets strong unforgeable security. The core is a branch structure that relies on secure cryptographic primitives. Inspired by the thought of conversion between backdoor and watermarking [14], we use the black-box undetectable backdoor to construct a black-box performance-loss watermarking for model.
However, a digital signature is an asymmetric cryptography technology, leading to its low computational efficiency. If Goldwasser et al.’s work [36] is used directly for model watermarking, the running overhead is too heavy to deploy in real scenarios. Therefore, we considered more efficient cryptographic techniques with similar security properties—collision-resistant hash functions and existentially unforgeable message authentication codes—and ultimately chose the latter to construct the watermarking.
III-C Planting a Parallel Branch in Model API
@@
Remote services have the advantages of a black box model, allowing watermarking modules to be added to the API. We propose a branch backdoor-based watermarking protocol for model. The protocol consists of two modules, the model module and the watermarking module, with the watermarking module further divided into the inner module and the outer module. The outer module includes a trigger generator and evidence verifier submodules. The inner module includes a trigger detector and evidence generator submodules deployed within the API.
III-C1 Overview
@@
Figure 2 shows the complete process of the branch watermarking protocol for the model. When an ownership dispute arises, the verifier uses the outer module to generate triggers, sends these triggers as query requests to the API, and verifies the returned responses to complete copyright certification. Specifically, when the API receives a query request, it forwards the input content to both the model module and the inner module (of the watermarking module), initially running in the service state. The model module performs the original inference process to obtain the output content.
@@
In the meantime, the inner module verifies if the current input is a trigger. If it is, the API transitions to the forensic state and proceeds to execute the evidence generator module. This module injects ownership information into the original output content of the model, obtaining ownership evidence. The API then returns this evidence as a response to the user. If the input is not a trigger, the API continues in the service state, returning the model’s original output as a response to the user. The triggers used in the forensic process are generated by the outer module. Importantly, the outer module is private or only open to a trusted third party.
@@
Faced with the two situations of dishonest behavior in the threat model, the owner of Model A can naturally use the outer module to generate triggers, conduct investigations on suspicious remote application services, and verify the evidence. In the first case, sending triggers to the suspicious application that pretends to use Model A would yield no ownership evidence. In the second case, sending triggers to an application claiming to use Model B but with performance closely resembling Model A would yield the ownership evidence of Model A.
III-C2 Defining Watermarking Protocol